OpenStudy (anonymous):

1) x^2+1/6x-9x^2 dx 2) xdx/2+6x+9x^2 3) 1/x^2-1 dx

OpenStudy (anonymous):

What are you supposed to to?

OpenStudy (anonymous):

Find antiderivative

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

An antiderivative of a function f is a function whose derivative is f . In other words, F is an antiderivative of f if F' = f . To find an antiderivative for a function f , we can often reverse the process of differentiation. For example, if f = x^4 , then an antiderivative of f is F = 1/5 x^5 , which can be found by reversing the power rule. Notice that not only is 1/5 x^5 an antiderivative of f , but so are 1/5 x^5 + 4 , 1/5 x^5 + 6 , etc. In fact, adding or subtracting any constant would be acceptable.

OpenStudy (anonymous):

That I get, but I just don't know how we antideriverate those functions

OpenStudy (anonymous):

Ok, First try to antideriverate them without the numbers.

OpenStudy (anonymous):

Also to do it part by part. So the first one do the numirator and then the denominater.

OpenStudy (anonymous):

Just do the same thing on all three.

OpenStudy (anonymous):

What, I don' get what u mean?

OpenStudy (anonymous):

For the first one do x^2+1 After that do 6x-9x^2

OpenStudy (anonymous):

For the second one do x Then do 2+6x+9x^2

OpenStudy (anonymous):

Then for the last one do 1first Then do x^2-1

OpenStudy (anonymous):

So... The first one will be \[\frac{ \frac{ x^3 }{ 3 }+x }{ 3x^2 -3x^3}\] The second one should be \[\frac{ \frac{x^2 }{ 2 } }{ 2x+3x^2 +3x^3}\] And the third one should be \[\frac{ x }{ \frac{ x^3 }{ 3 } -1 }\]

OpenStudy (irishboy123):

first one is \( \int \dfrac{x^2+1}{6x-9x^2} dx\) right?? if so you need to use some algebra, eg partial fractions, to simplify that and make it fit some known integration patterns i'd just pile in with with long division as this stuff can get really, really tedious after a while so you have \(\dfrac{x^2+1}{6x-9x^2} = \dfrac{1}{x}\dfrac{x^2+1}{6-9x}\) and why not try \((x^2 + 1) \div (-9x + 6)\) in ye olde long division format

OpenStudy (anonymous):

Oh lawd jesus. ain't nobody got time for that!

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

the first on i say

OpenStudy (irishboy123):

PS this is a really good resource http://www.mathsisfun.com/algebra/partial-fractions.html